# Timely Information Updating for Mobile Devices Without and With ML Advice

### v1

@@ Abstract
This paper investigates an information update system in which a mobile device monitors a physical process and sends status updates to an access point (AP). A fundamental trade-off arises between the timeliness of the information maintained at the AP and the update cost incurred at the device. To address this trade-off, we propose an online algorithm that determines when to transmit updates using only available observations. The proposed algorithm asymptotically achieves the optimal competitive ratio against an adversary that can simultaneously manipulate multiple sources of uncertainty, including the operation duration, the information staleness, the update cost, and the availability of update opportunities. Furthermore, by incorporating machine learning (ML) advice of unknown reliability into the design, we develop an ML-augmented algorithm that asymptotically attains the optimal consistency-robustness trade-off, even when the adversary can additionally corrupt the ML advice. The optimal competitive ratio scales linearly with the range of update costs, but is unaffected by other uncertainties. Moreover, an optimal competitive online algorithm exhibits a threshold-like response to the ML advice: it either fully trusts or completely ignores the ML advice, as partially trusting the advice cannot improve the consistency without severely degrading the robustness. Extensive simulations in stochastic settings further validate the theoretical findings in the adversarial environment.
@@ I Introduction
In recent years, the demand for timely information has surged across diverse systems. In Internet-of-Things (IoT) networks (e.g., unmanned aerial vehicles deployed for disaster response [gupta2015survey]), each IoT device is equipped with sensors (e.g., GPS, radar, and temperature sensors) that continuously monitor its surroundings. These sensors generate status updates about physical processes and transmit them to a central controller. By aggregating such updates, the controller has a real-time view of the environment, thereby enabling intelligent decision-making. Similarly, in location-based smartphone applications (e.g., navigation and gaming [karki2020characterizing]), users frequently report their locations to a central server so that the service can respond in real time. In both cases, a central entity relies on timely status updates from mobile devices to perform time-sensitive inference tasks.
To quantify the timeliness of information maintained at a central entity, Kaul, Yates, and Gruteser introduced the age of information metric in [kaul2012real], defined as the time elapsed since the most recently received update was generated. Under this definition, the information at the central entity linearly ages with time until it is updated. In addition to the linear aging function, more general nonlinear aging functions [kosta2017age] have also been analyzed. These functions further characterize the quality of an update, e.g., capturing how quickly the information held by the central entity deviates from the true status, or representing the penalty associated with using outdated information in decision-making. In this paper, we consider general aging functions.
@@
The scheduling problem is complicated by several forms of uncertainty in mobile networks: 1) The operation duration is uncertain, e.g., the runtime of a location-based application depends on how long a smartphone user keeps the application active. 2) The age increment may vary over time, e.g., a location update becomes stale more quickly when the device moves at a higher speed. 3) The update cost is also time-varying, e.g., user mobility causes fluctuations in energy consumption. 4) The device’s update opportunities may also be intermittent. Such cases include sporadic update arrivals (e.g., due to misalignment between the update generation periods and transmission slots) and transmission constraints that prevent the device from sending in certain slots (e.g., due to a power-saving policy [lin2022survey] or uplink scheduling decisions imposed by the AP [takeda2020understanding]).
Most prior works modeled uncertainties using stationary stochastic processes, e.g., employing an M/M/1 queueing model in [kaul2012real] to represent the update arrival and service processes. However, such assumptions are often unrealistic, e.g., when a device moves arbitrarily so that the service time no longer follows an exponential distribution. Even if such models fit reality reasonably well, the operation duration may be too short for the process to converge to stationarity. Moreover, practical stochastic models for operation duration or the information aging are often unclear. Such non-stationary uncertainties pose the central challenge for our scheduling design. Under non-stationarity, a scheduling algorithm cannot rely on future knowledge and must instead operate solely based on past and present observations, as an online algorithm.
@@
The above guarantee holds in the worst case (also referred to as the adversarial environment), across all possible uncertainty instances. However, such worst-case analysis can be overly pessimistic, since in practice future events often follow patterns that would be predicted using machine learning (ML). Motivated by this, Lykouris and Vassilvitskii [lykouris2018competitive] proposed incorporating (potentially imperfect) ML advice into online algorithms, as an approach that goes beyond the worst-case analysis. A central challenge in this setting is that the reliability of the ML advice is generally unknown. Thus, the design goal is twofold: 1) when the ML advice is accurate, the algorithm should perform well; 2) when the advice is unreliable, the algorithm should still provide performance guarantees. However, it is impossible to achieve both properties simultaneously. For example, an algorithm that blindly trusts the ML advice performs excellently under accurate ML advice but can suffer arbitrarily poor performance when the ML advice is wrong. Hence, the objective is to optimally balance these two properties.
Our second contribution is to integrate ML advice (specifically, advising the next update time) into our online scheduling framework. We introduce a hyperparameter $\lambda\in(0,1]$ to control the level of trust in the ML advice, where a smaller $\lambda$ places greater reliance on the ML advice. Our main result is that the proposed algorithm achieves the following asymptotic trade-off: 1) it achieves a total cost at most $\frac{\lambda e^{\lambda/R}}{e^{\lambda/R}-1}=\mathcal{O}(R)$ (known as the consistency) times the cost of blindly following the ML advice; 2) it also achieves a total cost at most $\frac{e^{\lambda/R}}{e^{\lambda/R}-1}=\mathcal{O}(\frac{R}{\lambda})$ (known as the robustness) times the minimum cost achieved by an optimal offline algorithm. Again, this balance depends only on the ratio $R$ and turns out to be optimal. We can observe that partially trusting the ML advice with $\lambda\in(0,1)$ cannot leverage the ML advice, as it yields (asymptotically) no improvement in the consistency. Thus, an optimal online algorithm almost displays a threshold-type behavior with respect to ML advice, either fully adopting it or ignoring it altogether.
@@ II Related Work
Extensive research has been conducted on analyzing and minimizing the age of information in diverse system settings. For example, Costa et al. [costa2016age] derived closed-form expressions for the average age in single-source systems; then, Yates and Kaul [yates2018age] extended the analysis to multi-source scenarios. Building on the foundational work, numerous system design strategies have been proposed to minimize the age, including scheduling algorithms [kadota2018scheduling], resource allocation schemes [park2020centralized], and sampling strategies [ornee2019sampling]. Beyond solely minimizing the age, several trade-off problems have also been investigated, such as age–throughput trade-off [mankar2021throughput, wang2025understanding] and age–energy trade-off [nath2018optimum, gu2019timely]. A comprehensive survey of these efforts is provided in [yates2021age].
Most prior age-related works assume stationary stochastic processes to model uncertainties (e.g., [yates2018age, kadota2018scheduling, park2020centralized, ornee2019sampling, mankar2021throughput, wang2025understanding, nath2018optimum, gu2019timely]). Since such assumptions can be overly optimistic, several works have examined how non-stationary (adversarial) environments affect information timeliness from different perspectives. Examples include adversarial ON/OFF channels [tseng2019online, sinha2022optimizing], adversarial update arrivals [saurav2023online], and adversarial aging functions [lin2025optimal, tripathi2021online]. Recent work [liu2025learning] further incorporated ML advice into online algorithms for adversarial ON/OFF channels.
@@ III-C Scheduling Classification
A scheduling algorithm is called a (cost-agnostic) online scheduling algorithm if it requires only the constants $C_{M}$ and $C_{m}$ and the realized age increments up to the current slot. For simplicity, we will omit the term cost-agnostic, with the understanding that all references to online scheduling algorithms refer to the cost-agnostic setting. Without access to the complete uncertainty instance, an online algorithm is generally unable to achieve the minimum total cost attainable by an optimal offline scheduling algorithm. Given an instance $\mathcal{I}$, let $\mathrm{OPT}(\mathcal{I})$ denote the minimum total cost achieved by an optimal offline algorithm. We evaluate the performance of online algorithms in terms of their competitiveness [buchbinder2009design] relative to the offline optimum, defined as follows.
@@
Note that in addition to the competitive ratio, another common performance metric is regret [lattimore2020bandit]. Regret measures the additive performance gap between an online learning approach and the best offline algorithm restricted to fixed decision rules. In contrast, the competitive ratio compares the performance of an online algorithm against the best offline algorithm without such restrictions. Moreover, online learning approaches typically provide regret guarantees only as the decision horizon grows to infinity. Such guarantees are less suitable for our setting, since a device may monitor and transmit updates only for a short and unpredictable duration.
@@
This paper considers the setting where the ML advice may be untrusted. While following perfect ML advice (which can minimize the total cost) yields the minimum total cost, blindly trusting imperfect ML advice (which cannot minimize the total cost) may lead to poor performance. Moreover, we assume that the reliability of the ML advice is unknown a priori. In this context, we characterize online scheduling algorithms with ML in terms of two metrics introduced in [lykouris2018competitive]: consistency and robustness. Consistency quantifies performance relative to the ML advice, while robustness guarantees performance in the worst case. These notions are formally defined as follows.
@@
IV Linear program formulation
The main challenge in designing our scheduling algorithm arises from several forms of uncertainty that are impractical to model using stationary stochastic processes and also limited current observations. To address these challenges, we use online algorithm design techniques based on linear programming [buchbinder2009design, bamas2020primal].
However, casting our problem as a linear program (LP) is non-trivial due to the non-linear nature of the age cost. For example, consider a scenario where the device transmits an update in slot 11 and schedules the next update in slot $x$. If the age increases linearly by one unit per slot until the next update, the cumulative age from slot 11 to slot $x$ is given by $\sum_{t=1}^{x}t=x(x+1)/2$, which grows quadratically with the decision variable $x$. See [arafa2017age] for a concrete example illustrating this behavior. This quadratic growth implies that the total age cost $\sum_{t=1}^{T}A(t)$ in Eq. (2) includes non-linear terms, thereby complicating direct LP formulation.
@@ IV-B LP formulation
We note that the clear/idle behavior in the virtual queueing system directly corresponds to sending/withholding an acknowledgment (ACK) to clear all received packets in the Transmission Control Protocol (TCP). From this perspective, we can leverage prior studies for the classic online TCP ACK problem [buchbinder2009design, Chapter 12] to formulate our offline optimal scheduling problem as an integer program with a linear objective and constraint functions:
@@
Before solving our online problem, we remark that, via the virtual queue transformation, our formulation generalizes the classical online TCP ACK problem [buchbinder2009design, Chapter 12] and its learning-augmented variant [bamas2020primal]. In the classical TCP ACK setting, each ACK incurs a constant cost. In contrast, our objective in Eq. (4a) allows the ACK (i.e., clearing) cost to vary across slots. Later, in Section VII, we further generalize the problem to scenarios where the ACK channel alternates between ON and OFF states, and transmitting an ACK during an ON slot incurs an adversarially chosen cost. This generalization is practically relevant in noisy wireless environments; moreover, it poses theoretical challenges, since the adversary simultaneously controls multiple sources of uncertainty and an online algorithm is restricted to transmitting an ACK only in ON slots. As we will show, our generalized setting also yields a fundamentally different optimal competitive ratio that depends solely on the cost range (i.e., time-varying costs are the dominant factor). When augmented with ML advice, our generalized setting exhibits behavior that differs qualitatively from the classic learning-augmented TCP model; in particular, it yields a threshold-like optimal trust rule when the cost range is large.
@@ V-A Online LP algorithm
The underlying idea is that in each slot $t$, our scheduling algorithm that will be proposed in Section V-C makes a probabilistic decision: to set $d(t)=1$ with some probability or $d(t)=0$ otherwise. The probability is governed by the current value of $x(t)$, which is determined in Line 1 of Alg. 1. Accordingly, $x(t)$ can be interpreted as the probability of clearing the virtual queue in slot $t$. In this context, the cumulative sum $\sum_{\tau=T_{i}}^{t}x(\tau)$ represents the cumulative clearing probability (up to slot $t$) for virtual packet ii.
@@ 11    $x(t)\leftarrow x(t)+\frac{1}{C_{M}}\sum_{\tau=T_{i}}^{t}x(\tau)+\frac{1}{\theta C_{M}}$;
In this section, we analyze the objective value in Eq. (5a) computed by Alg. 1. Unlike prior studies [buchbinder2009design, bamas2020primal] that analyze online algorithms for LPs using primal–dual techniques, our analysis exploits structural properties of Alg. 1 and its relation to an optimal offline scheduling algorithm. An advantage of our approach is that it provides a unified analysis framework for all proposed LP algorithms (including Algs. 1, 3, and 4) without the need to construct separate dual solutions for different scenarios.
@@ 23 $d(t)\leftarrow 0$;
Leveraging the fractional-to-probabilistic conversion technique proposed in [buchbinder2009design, Chapter 12], this section presents a randomized online scheduling algorithm in Alg. 2, which converts the fractional solution $x(t)$ generated by Alg. 1 into a probabilistic transmission decision.
@@
Fix an instance $\mathcal{I}$. Following [tseng2019online], we can show that the expected clearing cost in each slot $t$ under Alg. 2 is upper bounded by the value of $C(t)x(t)$ as computed by Alg. 1. Similarly, the expected number of virtual packets present in slot $t$ under Alg. 2 is upper bounded by $z_{i}(t)$ as computed by Alg. 1. Therefore, the expected total cost in Eq. (3) incurred by Alg. 2 is bounded above by the objective value in Eq. (5a) computed by Alg. 1. The result then follows directly from Theorem 6. ∎
@@
(Sketch) Consider the initial age $A_{0}=1$, a fixed age increment sequence $\boldsymbol{\Delta A}=(0,\cdots,0)$, and a fixed update cost sequence $\mathbf{C}=(C_{m},C_{m}R,C_{m}R,\cdots,C_{m}R)$. Only the operation duration $T$ is unknown to the device. The age increment sequence models a scenario where information becomes stale very slowly or no more aging when there is no content change in the process the device is monitoring as in [salimnejad2025age]. Despite this single source of unknown uncertainty, we show in Appendix D that as $C_{m}\to\infty$, no online scheduling algorithm can achieve a competitive ratio smaller than $(e^{1/R})/(e^{1/R}-1)$. ∎
Through the lower bound on the competitive ratio in Lemma 9 and the matching achievability scheme proposed in Alg. 2, we establish that the optimal competitive ratio against an adversary that can jointly manipulate the operation duration, the age increment, and the update cost is $(e^{1/R})/(e^{1/R}-1)$. This matches the result in the classic online TCP [buchbinder2009design] for $R=1$ (without cost variation). Moreover, this ratio is $\mathcal{O}(R)$ for large $R$, scaling linearly with the update cost range $R$, while is unaffected by all other sources of uncertainty. Thus, when the cost fluctuates, it becomes fundamentally harder for any online algorithm to balance timeliness and update cost. Moreover, see Fig. 3 for the competitive ratio at finite values of $R$, which also appears approximately linear in $R$.
@@
This remark discusses how Alg. 1 can be adapted to scenarios in which the bounds $C_{M}$ and $C_{m}$ on the update cost are not known in advance. In this case, we propose to periodically estimate the update cost using channel estimation techniques (e.g., see [liu2025learning]). Let $T_{\text{est}}$ denote the estimation period and assume that the update cost $C(t)$ is measured in slots $t=1,\,T_{\text{est}}+1,\,2T_{\text{est}}+1,\,\cdots$. For each slot $t$, by $C_{M}(t)=\max_{0\leq n\leq(t-1)/T_{\text{est}}}C(nT_{\text{est}}+1)$ and $C_{m}(t)=\min_{0\leq n\leq(t-1)/T_{\text{est}}}C(nT_{\text{est}}+1)$ we define the maximum and minimum observed costs up to slot $t$, respectively. In Alg. 2 (and the corresponding Alg. 1), we replace $C_{M}$ and $C_{m}$ by $C_{M}(t)$ and $C_{m}(t)$, respectively. This remains an online algorithm, as no future information is used. Let $\theta(t)=(1+(1/(C_{M}(t)))^{C_{m}(t)}-1$. Then, the increment of the objective value in Eq. (7) becomes
@@
As $C_{u}\to\infty$, the modified online algorithm achieves the asymptotic competitive ratio of $\mathcal{O}(e^{1/R}/(e^{1/R}-1))$, which matches the order of the original competitive ratio (but with an inflated multiplicative pre-constant $1+(T_{\text{est}}-1)\Delta\mathcal{E}_{M}/\mathcal{E}_{m}$). The same fix can apply to all remaining algorithms proposed later, yielding the same pre-constant.

### v2

@@ Abstract
This paper investigates an information update system in which a mobile device monitors a physical process and sends status updates to an access point (AP). A fundamental trade-off arises between the timeliness of the information maintained at the AP and the update cost incurred at the device. To address this trade-off, we propose an online algorithm that determines when to transmit updates using only available observations. The proposed algorithm asymptotically achieves the optimal competitive ratio against an adversary that can simultaneously manipulate multiple sources of uncertainty, including the operation duration, information staleness, update cost, and update opportunities. Furthermore, by incorporating machine learning (ML) advice of unknown reliability into the design, we develop an ML-augmented algorithm that asymptotically attains the optimal consistency-robustness trade-off, even when the adversary can additionally corrupt the ML advice. The optimal competitive ratio scales linearly with the range of update costs, but is unaffected by other sources of uncertainty. Moreover, an optimal competitive online algorithm exhibits a threshold-like response to the ML advice: it either fully trusts or completely ignores the ML advice, as partially trusting the advice cannot improve the consistency without severely degrading the robustness. Extensive simulations in stochastic settings further validate the theoretical findings in the adversarial environment.
@@ I Introduction
In recent years, the demand for timely information has surged across diverse systems. In Internet-of-Things (IoT) networks (e.g., unmanned aerial vehicles deployed for disaster response [7]), each IoT device is equipped with sensors (e.g., GPS, radar, and temperature sensors) that continuously monitor its surroundings. These sensors generate status updates about physical processes and transmit them to a central controller. By aggregating such updates, the controller has a real-time view of the environment, thereby enabling intelligent decision-making. Similarly, in location-based smartphone applications (e.g., navigation and gaming [10]), users frequently report their locations to a central server so that the service can respond in real time. In both cases, a central entity relies on timely status updates from mobile devices to perform time-sensitive inference tasks.
To quantify the timeliness of information maintained at a central entity, Kaul, Yates, and Gruteser introduced the age of information metric in [11], defined as the time elapsed since the most recently received update was generated. Under this definition, the information at the central entity linearly ages with time until it is updated. In addition to the linear aging function, more general nonlinear aging functions [12] have also been analyzed. These functions further characterize the quality of an update, e.g., capturing how quickly the information held by the central entity deviates from the true status, or representing the penalty associated with using outdated information in decision-making. In this paper, we consider general aging functions.
@@
The scheduling problem is complicated by several forms of uncertainty in mobile networks: 1) The operation duration is uncertain, e.g., the runtime of a location-based application depends on how long a smartphone user keeps the application active. 2) The age increment may vary over time, e.g., a location update becomes stale more quickly when the device moves at a higher speed. 3) The update cost is also time-varying, e.g., user mobility causes fluctuations in energy consumption. 4) The device’s update opportunities may also be intermittent. Such cases include sporadic update arrivals (e.g., due to misalignment between the update generation periods and transmission slots) and transmission constraints that prevent the device from sending in certain slots (e.g., due to a power-saving policy [14] or uplink scheduling decisions imposed by the AP [24]).
Most prior works modeled uncertainties using stationary stochastic processes, e.g., employing an M/M/1 queueing model in [11] to represent the update arrival and service processes. However, such assumptions are often unrealistic, e.g., when a device moves arbitrarily so that the service time no longer follows an exponential distribution. Even if such models fit reality reasonably well, the operation duration may be too short for the process to converge to stationarity. Moreover, practical stochastic models for operation duration or the information aging are often unclear. Such non-stationary uncertainties pose the central challenge for our scheduling design. Under non-stationarity, a scheduling algorithm cannot rely on future knowledge and must instead operate solely based on past and present observations, as an online algorithm.
@@
The above guarantee holds uniformly over all uncertainty instances, including the worst-case scenario. Such worst-case analysis can be overly pessimistic, since in practice future events often follow patterns that would be predicted using machine learning (ML). Motivated by this, Lykouris and Vassilvitskii [17] proposed incorporating (potentially imperfect) ML advice into online algorithms, as an approach that goes beyond the worst-case analysis. A central challenge in this setting is that the reliability of the ML advice is generally unknown. Thus, the design goal is twofold: 1) when the ML advice is accurate, the algorithm should perform well; 2) when the advice is unreliable, the algorithm should still provide performance guarantees. However, it is impossible to achieve both properties simultaneously. For example, an algorithm that blindly trusts the ML advice performs excellently under accurate ML advice but can suffer arbitrarily poor performance when the ML advice is wrong. Hence, the objective is to optimally balance these two properties.
Our second contribution is to integrate untrusted ML advice (specifically, advice on the next update time) into our online scheduling framework. We introduce a hyperparameter $\lambda\in(0,1]$ to control the level of trust in the ML advice, where a smaller $\lambda$ places greater reliance on the ML advice. Our main result is that the proposed algorithm achieves the following asymptotic trade-off: 1) it achieves a total cost at most $\frac{\lambda e^{\lambda/R}}{e^{\lambda/R}-1}=\mathcal{O}(R)$ (known as the consistency) times the cost of blindly following the ML advice; 2) it also achieves a total cost at most $\frac{e^{\lambda/R}}{e^{\lambda/R}-1}=\mathcal{O}(\frac{R}{\lambda})$ (known as the robustness) times the minimum cost achieved by an optimal offline algorithm. Again, this balance depends only on the ratio $R$ and turns out to be optimal. We can observe that partially trusting the ML advice with $\lambda\in(0,1)$ cannot leverage the ML advice, as it yields (asymptotically) no improvement in the consistency. Thus, an optimal online algorithm displays an essentially threshold-type behavior with respect to ML advice, either fully adopting it or ignoring it altogether.
@@ II Related work
Extensive research has been conducted on analyzing and minimizing the age of information in diverse system settings. For example, Costa et al. [5] derived closed-form expressions for the average age in single-source systems; then, Yates and Kaul [28] extended the analysis to multi-source scenarios. Building on the foundational work, numerous system design strategies have been proposed to minimize the age, including scheduling algorithms [9], resource allocation schemes [21], and sampling strategies [20]. Beyond solely minimizing the age, several trade-off problems have also been investigated, such as age–throughput trade-off [18, 27] and age–energy trade-off [19, 6]. A comprehensive survey of these efforts is provided in [29].
Most prior age-related works assume stationary stochastic processes to model uncertainties (e.g., [28, 9, 21, 20, 18, 27, 19, 6]). Since such assumptions can be overly optimistic, several works have examined how non-stationary (adversarial) environments affect information timeliness from different perspectives. Examples include adversarial ON/OFF channels [26, 23], adversarial update arrivals [22], and adversarial aging functions [15, 25]. Recent work [16] further incorporated ML advice into online algorithms for adversarial ON/OFF channels.
@@ III-C Scheduling algorithm classification
A scheduling algorithm is called a (cost-agnostic) online scheduling algorithm if it requires only the constants $C_{M}$ and $C_{m}$ and the realized age increments up to the current slot. For simplicity, we will omit the term cost-agnostic, with the understanding that all references to online scheduling algorithms refer to the cost-agnostic setting. Without access to the complete uncertainty instance, an online algorithm is generally unable to achieve the minimum total cost attainable by an optimal offline scheduling algorithm. Given an instance $\mathcal{I}$, let $\mathrm{OPT}(\mathcal{I})$ denote the minimum total cost achieved by an optimal offline algorithm. We evaluate the performance of online algorithms in terms of their competitiveness [4] relative to the offline optimum, defined as follows.
@@
Note that in addition to the competitive ratio, another common performance metric is regret [13]. Regret measures the additive performance gap between an online learning approach and the best offline algorithm restricted to fixed decision rules. In contrast, the competitive ratio compares the performance of an online algorithm against the best offline algorithm without such restrictions. Moreover, online learning approaches typically provide regret guarantees only as the decision horizon grows to infinity. Such guarantees are less suitable for our setting, since a device may monitor and transmit updates only for a short and unpredictable duration.
@@
This paper considers the setting where the ML advice may be untrusted. While following perfect ML advice (which can minimize the total cost) yields the minimum total cost, blindly trusting imperfect ML advice (which cannot minimize the total cost) may lead to poor performance. Moreover, we assume that the reliability of the ML advice is unknown a priori. In this context, we characterize online scheduling algorithms with ML in terms of two metrics introduced in [17]: consistency and robustness. Consistency quantifies performance relative to the ML advice, while robustness guarantees performance in the worst case. These notions are formally defined as follows.
@@
IV Linear program formulation for offline scheduling
The main challenge in designing our scheduling algorithm arises from several forms of uncertainty that are impractical to model using stationary stochastic processes and also limited current observations. To address these challenges, we leverage online algorithm design techniques based on offline optimal linear programming [4, 3].
However, casting our offline scheduling problem as an linear program (LP) is non-trivial due to the non-linear nature of the age cost. For example, consider a scenario where the device transmits an update in slot $1$ and schedules the next update in slot $x$. If the age increases linearly by one unit per slot until the next update, the cumulative age from slot $1$ to slot $x$ is given by $\sum_{t=1}^{x}t=x(x+1)/2$, which grows quadratically with the decision variable $x$. See [2] for a concrete example illustrating this behavior. This quadratic growth implies that the total age cost $\sum_{t=1}^{T}A(t)$ in Eq. (2) includes non-linear terms, thereby complicating direct LP formulation.
@@ IV-B LP formulation
We note that the clear/idle behavior in the virtual queueing system directly corresponds to sending/withholding an acknowledgment (ACK) to clear all received packets in the Transmission Control Protocol (TCP). From this perspective, we can leverage prior studies for the classic online TCP ACK problem [4, Chapter 12] to formulate our offline optimal scheduling problem as an integer program with a linear objective and constraint functions:
@@
Before solving our online problem, we remark that, via the virtual queue transformation, our formulation generalizes the classical online TCP ACK problem [4, Chapter 12] and its learning-augmented variant [3]. In the classical TCP ACK setting, each ACK incurs a constant cost. In contrast, our objective in Eq. (4a) allows the ACK (i.e., clearing) cost to vary across slots. Later, in Section VII, we further generalize the problem to scenarios where the ACK channel alternates between ON and OFF states, and transmitting an ACK during an ON slot incurs an adversarially chosen cost. This generalization is practically relevant in noisy wireless environments; moreover, it poses theoretical challenges, since the adversary simultaneously controls multiple sources of uncertainty and an online algorithm is restricted to transmitting an ACK only in ON slots. As we will show, our generalized setting also yields a fundamentally different optimal competitive ratio that depends solely on the cost range (i.e., time-varying costs are the dominant factor). When augmented with ML advice, our generalized setting exhibits behavior that differs qualitatively from the classic learning-augmented TCP model; in particular, it yields a threshold-like optimal trust rule when the cost range is large.
@@ V-A Online LP algorithm
The underlying idea is that in each slot $t$, our scheduling algorithm that will be proposed in Section V-C makes a probabilistic decision: to set $d(t)=1$ with some probability or $d(t)=0$ otherwise. The probability is governed by the current value of $x(t)$, which is determined in Line 1 of Alg. 1. Accordingly, $x(t)$ can be interpreted as the probability of clearing the virtual queue in slot $t$. In this context, the cumulative sum $\sum_{\tau=T_{i}}^{t}x(\tau)$ represents the cumulative clearing probability (up to slot $t$) for virtual packet $i$.
@@ 11    $x(t)\leftarrow x(t)+\frac{1}{C_{M}}\sum_{\tau=T_{i}}^{t}x(\tau)+\frac{1}{\theta C_{M}}$;
In this section, we analyze the objective value in Eq. (5a) computed by Alg. 1. Unlike prior studies [4, 3] that analyze online algorithms for LPs using primal–dual techniques, our analysis exploits structural properties of Alg. 1 and its relation to an optimal offline scheduling algorithm. An advantage of our approach is that it provides a unified analysis framework for all proposed LP algorithms (including Algs. 1, 3, and 4) without the need to construct separate dual solutions for different scenarios.
@@ 23 $d(t)\leftarrow 0$;
Leveraging the fractional-to-probabilistic conversion technique proposed in [4, Chapter 12], this section presents a randomized online scheduling algorithm in Alg. 2, which converts the fractional solution $x(t)$ generated by Alg. 1 into a probabilistic transmission decision.
@@
Fix an instance $\mathcal{I}$. Following [26], we can show that the expected clearing cost in each slot $t$ under Alg. 2 is upper bounded by the value of $C(t)x(t)$ as computed by Alg. 1. Similarly, the expected number of virtual packets present in slot $t$ under Alg. 2 is upper bounded by $z_{i}(t)$ as computed by Alg. 1. Therefore, the expected total cost in Eq. (3) incurred by Alg. 2 is bounded above by the objective value in Eq. (5a) computed by Alg. 1. The result then follows directly from Theorem 6. ∎
@@
(Sketch) Consider the initial age $A_{0}=1$, a fixed age increment sequence $\boldsymbol{\Delta A}=(0,\cdots,0)$, and a fixed update cost sequence $\mathbf{C}=(C_{m},C_{m}R,C_{m}R,\cdots,C_{m}R)$. Only the operation duration $T$ is unknown to the device. The age increment sequence captures a versioned monitoring scenario, where the age remains unchanged until a new version is generated, as in the age of version metric [1]. Despite this single source of unknown uncertainty, we show in Appendix D that as $C_{m}\to\infty$, no online scheduling algorithm can achieve a competitive ratio smaller than $(e^{1/R})/(e^{1/R}-1)$. ∎
Through the lower bound on the competitive ratio in Lemma 9 and the matching achievability scheme proposed in Alg. 2, we establish that the optimal competitive ratio against an adversary that can jointly manipulate the operation duration, the age increment, and the update cost is $(e^{1/R})/(e^{1/R}-1)$. This matches the result in the classic online TCP [4] for $R=1$ (without cost variation). Moreover, this ratio is $\mathcal{O}(R)$ for large $R$, scaling linearly with the update cost range $R$, while is unaffected by all other sources of uncertainty. Thus, when the cost fluctuates, it becomes fundamentally harder for any online algorithm to balance timeliness and update cost. Moreover, see Fig. 3 for the competitive ratio at finite values of $R$, which also appears approximately linear in $R$.
@@
This remark discusses how Alg. 1 can be adapted to scenarios in which the bounds $C_{M}$ and $C_{m}$ on the update cost are not known in advance. In this case, we propose to periodically estimate the update cost using channel estimation techniques (e.g., see [16]). Let $T_{\text{est}}$ denote the estimation period and assume that the update cost $C(t)$ is measured in slots $t=1,\,T_{\text{est}}+1,\,2T_{\text{est}}+1,\,\cdots$. For each slot $t$, by $C_{M}(t)=\max_{0\leq n\leq(t-1)/T_{\text{est}}}C(nT_{\text{est}}+1)$ and $C_{m}(t)=\min_{0\leq n\leq(t-1)/T_{\text{est}}}C(nT_{\text{est}}+1)$ we define the maximum and minimum observed costs up to slot $t$, respectively. In Alg. 2 (and the corresponding Alg. 1), we replace $C_{M}$ and $C_{m}$ by $C_{M}(t)$ and $C_{m}(t)$, respectively. This remains an online algorithm, as no future information is used. Let $\theta(t)=(1+(1/(C_{M}(t)))^{C_{m}(t)}-1$. Then, the increment of the objective value in Eq. (7) becomes
@@
As $C_{u}\to\infty$, the modified online algorithm achieves the asymptotic competitive ratio of $\mathcal{O}(e^{1/R}/(e^{1/R}-1))$, which matches the order of the original competitive ratio (but with an inflated multiplicative pre-constant $1+(T_{\text{est}}-1)(\Delta\mathcal{E}_{M}/\mathcal{E}_{m}$). The same fix can apply to all remaining algorithms proposed later, yielding the same pre-constant.